OpenStudy (anonymous):

Applicable to this and all other questions-- For any solution that involves two answers, type one answer separated by a comma and then the other answer. Solve the equation x2 = 81 by finding square roots.

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

x^2 = 81 you actually mean this sqrt(x^2)=sqrt(81) take sqrt of both sides x= 9 or x=-9 both work but since it says seperate answer by commas it is 9,-9

OpenStudy (anonymous):

x^2 = 81 you actually mean this sqrt(x^2)=sqrt(81) take sqrt of both sides x= 9 or x=-9 both work but since it says seperate answer by commas it is 9,-9

OpenStudy (anonymous):

omg thank you

OpenStudy (anonymous):

u r welcome

OpenStudy (anonymous):

hey she copied me :P

OpenStudy (anonymous):

hahahaha its ok lol

OpenStudy (anonymous):

no i looked it up in my mathematical term book ok so stop hatin

OpenStudy (anonymous):

lol one of those sounds nice lol